Prat Corp. started the 2018 accounting period with $32,000 of assets (all cash), $13,000 of liabilities, and $17,000 of common stock. During the year, the Retained Earnings account increased by $17,050. The bookkeeper reported that Prat paid cash expenses of $32,000 and paid a $3,200 cash dividend to the stockholders, but she could not find a record of the amount of cash that Prat received for performing services. Prat also paid $6,500 cash to reduce the liability owed to the bank, and the business acquired $7,200 of additional cash from the issue of common stock. a-1. Prepare an income statement for the 2018 accounting period. a-2. Prepare a statement of changes in stockholders’ equity for the 2018 accounting period. a-3. Prepare a period-end balance sheet for the 2018 accounting period. a-4. Prepare a statement of cash flows for the 2018 accounting period. B.) Determine the percentage of total assets that were provided by creditors, investors, and earnings.
Answer #1

Answer a-1:

Given:

Retained Earnings account increased by $17,050

Prat paid cash expenses of $32,000 and paid a $3,200 cash dividend to the stockholders.

Hence:

Net income = Increase in Retained Earnings + dividends paid = 17050 + 3200 = $20,250

Service revenue = Net income + expenses = 20250 + 32000 = $52,250

Answer a-2:

Beginning balance of retained earnings = Total assets - liabilities - common stock = 32000 - 13000 - 17000 = $2,000

Answer a-3:

Answer a-4:

Answer B:

Percentage of total assets that were provided by creditors = 6500/49750 = 13.07%

Percentage of total assets that were provided by investors = 24200/49750 = 48.64%

percentage of total assets that were provided by earnings.= 19050 /49750 = 38.29%

Hence:

Percentage of total assets that were provided by creditors =13.07%

Percentage of total assets that were provided by investors = 48.64%

percentage of total assets that were provided by earnings.= 38.29%